OpenStudy (anonymous):

x^2-49/4x+24*x+6/x-7

OpenStudy (anonymous):

Factor each of the four expressions where possible, x^2 - 49 = (x+7)(x-7) 4x + 24 = 4(x+6) x+6 cannot be factored further x-7 cannot be factored further So you have (x-7)(x+7)/4(x+6) times (x+6)/(x-7) The (x+6)'s cancels, the (x-7)'s cancels So we are left with (x+7)/4